I need to sketch a graph of the function:
f(x)= x+1, for x<2
2-x, for x>(or equal to)2
I can't figure out how to enter this into my T84 in order to graph. I recieved -1 of 4 poss. pts on my test so I know I am sort of on the right track.
Thank you!

Answer by jim_thompson5910(35256)

Graph the two equations y = x+1 and y = 2-x both on the same coordinate system. Make them different colors or shades so you can differentiate the two.

Now focus on y = x+1. Erase the portion that's to the right of x = 2 on this graph and draw an open circle at the endpoint

Then focus on y = 2-x. Erase the portion that's to the left of x = 2 on this graph and draw a closed circle at the endpoint